# CASteer: Cross-Attention Steering for Controllable Concept Erasure

### v3

@@ 1 Introduction
Recent advances in diffusion models DBLP:conf/nips/HoJA20; DBLP:conf/cvpr/RombachBLEO22 have revolutionized image DBLP:conf/iclr/PodellELBDMPR24 and video generation emu-video, achieving unprecedented realism. These models operate by gradually adding noise to data during a forward process and then learning to reverse this noise through a series of iterative steps, reconstructing the original data from randomness. By leveraging this denoising process, diffusion models generate high-quality, realistic outputs, making them a powerful tool for creative and generative tasks.
@@
Existing methods for concept erasure in diffusion models remain narrow in scope. LoRA-based fine-tuning DBLP:conf/iclr/HuSWALWWC22 is effective for removing specific objects or styles but struggles with abstract or composite concepts (e.g., nudity, violence, or ideological symbolism), and scales poorly when multiple concepts must be removed, requiring separate adapters or costly retraining. Prompt-based interventions DBLP:journals/corr/abs-2410-12761 offer greater flexibility for abstract harm reduction but lack precision in suppressing concrete attributes, often failing to generalize across concept variations. As a result, existing strategies fall short of delivering reliable, efficient, and broad-spectrum concept erasure.
In this work, we introduce CASteer, a training-free method for controllable concept erasure that leverages the principle of steering to influence hidden representations of diffusion models dynamically. Our method builds on recent findings that deep neural networks encode features into approximately linear subspaces elhage2021mathematical; DBLP:conf/nips/WuGIPG23. Prior research has shown that intermediate subspaces of diffusion backbones also exhibit this property, with directions that modulate the strength of particular features DBLP:conf/iclr/KwonJU23; DBLP:conf/nips/ParkKCJU23; DBLP:conf/cvpr/SiH0024; DBLP:conf/cvpr/TumanyanGBD23; DBLP:conf/cvpr/0010S00G24. Yet, these techniques remain limited in scope, often restricted to specific subspaces, requiring training, or offering only coarse control.
Our approach departs from this paradigm. We show that multiple subspaces within diffusion models exhibit linear properties that can be harnessed for precise concept erasure. For each concept of interest, we generate $k$ positive images (where $k\geq 1$) containing the concept and $k$ negative images not containing it, and compute the steering vectors by subtracting the averaged hidden representations of the network across negative images from those of positive ones. During inference, these precomputed vectors are applied directly to the model activations, allowing us to selectively suppress undesirable concepts without retraining or degrading the overall image quality. Experiments demonstrate that CASteer achieves fine-grained erasure of harmful or unwanted concepts (e.g., nudity, violence), while maintaining robustness across a wide range of diffusion models, including SD 1.4, SDXL DBLP:conf/iclr/PodellELBDMPR24, Sana xie2025sana, and their distilled variants (e.g., SDXL-Turbo DBLP:conf/eccv/SauerLBR24, Sana-Sprint chen2025sana-sprint).
@@ 2 Related work
Data-driven AI Safety. Ensuring the safety of image and text-to-image generative models hinges on preventing the generation of harmful or unwanted content. Common approaches include curating training data with licensed material adobefirefly; DBLP:conf/nips/SchuhmannBVGWCC22, fine-tuning models to suppress harmful outputs DBLP:conf/cvpr/RombachBLEO22; DBLP:journals/corr/abs-2006-11807, or deploying post-hoc content detectors nudenet; DBLP:journals/corr/abs-2210-04610. While promising, these strategies face critical limitations: data filtering introduces inherent biases DBLP:journals/corr/abs-2006-11807, detectors are computationally efficient but often inaccurate or easily bypassed DBLP:conf/iccv/GandikotaMFB23; remove_nsfw_detection, and model retraining becomes costly when new harmful concepts emerge. Alternative methods leverage text-domain interventions, such as prompt engineering DBLP:journals/corr/abs-2006-11807 or negative prompts DBLP:journals/corr/abs-2305-16807; DBLP:conf/cvpr/SchramowskiBDK23. Yet these remain vulnerable to adversarial attacks, lack flexibility and precision as they operate in the discrete space of tokens, and often fail to address the disconnect between text prompts and visual outputs—models can still generate undesired content even when text guidance is “safe.” Our approach instead operates in the joint image-text latent space of diffusion models, enabling more robust and granular control over generated content without relying solely on textual constraints.
@@
Model-driven AI Safety. Current methods DBLP:conf/iccv/GandikotaMFB23; DBLP:conf/iccv/KumariZWS0Z23; DBLP:conf/nips/HengS23; zhang2023forgetmenot; huang2023receler; lee2024cpe erase unwanted concepts by fine-tuning or otherwise optimising models and adapters to shift probability distributions toward null or surrogate tokens, often combined with regularization or generative replay DBLP:conf/nips/ShinLKK17. Other methods, such as DBLP:conf/wacv/GandikotaOBMB24; DBLP:conf/eccv/GongCWCJ24, use direct weight editing to remove unwanted concepts. Although effective, these approaches lack precision, inadvertently altering or removing unrelated concepts. Advanced techniques like SPM DBLP:conf/cvpr/Lyu0HCJ00HD24 and MACE DBLP:conf/cvpr/LuWLLK24 improve specificity through LoRA adapters DBLP:conf/iclr/HuSWALWWC22, transport mechanisms, or prompt-guided projections to preserve model integrity. However, while promising for concrete concepts (e.g., Mickey Mouse), they still struggle with abstract concepts (e.g., nudity) and require parameter updates. Another group of methods focuses on interventions into internal mechanisms of generative models. Methods like Prompt-to-Prompt DBLP:conf/iclr/HertzMTAPC23 enable fine-grained control over text-specified concepts (e.g., amplifying or replacing elements) through interventions to cross-attention maps, yet fail to fully suppress undesired content, particularly when concepts are implicit or absent from prompts. This task-specific specialization limits their utility for safety-critical erasure, where complete removal is required. CASteer bridges this gap, enabling precise, universal concept suppression without relying on textual priors or compromising unrelated model capabilities. Another area of research focuses on removing information about undesired concepts from text embeddings that generative models are conditioned on DBLP:journals/corr/abs-2410-12761; DBLP:journals/corr/abs-2410-02710; DBLP:journals/corr/abs-2411-10329. However, as these methods operate on a discrete space of token embeddings, their trade-off between the effectiveness of erasure and the preservation of other features is limited. zhang2024defensive proposes using adversarial training for concept unlearning; however, training this method is computationally intensive. In contrast, CASteer eliminates training entirely, enabling direct, non-invasive concept suppression in the model’s latent space without collateral damage to unrelated features.
@@
Utilizing directions in latent spaces. This area of research focuses on finding interpretable directions in various intermediate spaces of diffusion models DBLP:conf/iclr/KwonJU23; DBLP:conf/nips/ParkKCJU23; DBLP:conf/cvpr/SiH0024; DBLP:conf/cvpr/TumanyanGBD23, which can then be used to control the semantics of generated images. Based on this idea, SDID DBLP:conf/cvpr/0010S00G24 recently proposed to learn a vector for each given concept, which is then added to the intermediate activation of a bottleneck layer of the diffusion model during inference to provoke the presence of this concept in the generated image. However, this method is highly architecture-specific and fails to deliver precise control over attributes. In our work, we propose a training-free method for constructing interpretable directions in intermediate activation spaces of various diffusion models for more precise control of image generation. SAeUron cywinski2025saeuron utilizes Sparse Autoencoders Olshausen1997SparseCW (SAEs) to find interpretable directions in the activation space of the diffusion model. However, SAEs are unstable, require extensive training, and do not provide initial control over the set of attributes that can be erased. In contrast, CASteer does not require training and provides direct control over the manipulated attributes.
@@ 3.1 Choice of layers to steer
Most modern diffusion models use U-Net or Diffusion Transformers (DiT) DBLP:conf/iccv/PeeblesX23 as a backbone. They consist of a set of Transformer blocks, each having three main components: cross-attention (CA) layer, self-attention (SA) layer, and MLP layer, all of which contribute to the residual stream of the model. Among those, CA layers are the only place in the model where information from the text prompt goes into the model, guiding text-to-image generation. For every image patch and prompt embedding, each CA layer generates a vector matching the size of the image patch embedding. After summation, these vectors transmit text-prompt information to corresponding image regions DBLP:conf/iclr/HertzMTAPC23.
As the semantics of the resulting image is mostly determined by the text prompt, we modify the outputs of the CA layers during inference, which results in effective, yet precise, control over the features of the generated image. Thus, CASteer constructs steering vectors for the outputs of every CA layer in the model. In the supplementary, we present experiments on applying CASteer to steer outputs of other layers (SA, MLP, and outputs of intermediate layers inside CA blocks).
@@ 3.2 Construction of steering vectors
Figure 1: Main pipeline. (Bottom left, gray background) For computing a steering vector, we prompt diffusion model with two prompts that differ in a desired concept, e.g., “anime style” and save CA outputs at each timestamp $t$ and each CA layer $i$. We average these outputs over image patches and get averaged CA outputs $ca^{pos\_avg}_{it}$ and $ca^{neg\_avg}_{it}$ for each $t$ and $i$. We subtract the latter from the former, getting a steering vector for the layer $i$ and timestamp $t$ $ca^{anime}_{it}$. (Right) For deleting concept $X$ from generation, at each denoising step $t$, we subtract steering vector $ca^{X}_{it}$ multiplied by intensity $\alpha$ from the CA outputs of the layer $i$.
@@ 3.3 Using steering vectors to control generation
We propose to estimate $\alpha$ for concept deletion by using the dot product between $ca^{X}_{it}$ and corresponding CA output $ca^{out}_{itk}$ ($\langle ca^{X}_{it},ca^{out}_{itk}\rangle$) as an assessment of amount of $X$ that is present in the image part corresponding to $k^{th}$ patch of $ca^{out}_{it}$. As $ca^{X}_{it}$ is normalized, the value of this dot product is the length of the projection of the CA output $ca^{out}_{itk}$ onto the steering vector $ca^{X}_{it}$. As $ca^{X}_{it}$ can be seen as a direction in a linear subspace corresponding to the concept $X$, the length of the projection can be seen as the amount of $X$ that is present in $ca^{out}_{itk}$. That said, for removing information about $X$ from $ca^{out}_{itk}$, we propose to subtract the amount of $ca^{X}_{it}$ proportionate to the dot product between $ca^{X}_{it}$ and $ca^{out}_{itk}$ from $ca^{out\_new}_{itk}$, i.e., define $\alpha=\beta(ca^{X}_{it},ca^{out}_{itk})$. Consequently, Eq. 3 becomes the following:
@@ 3.4 Practical considerations
Steering multiple concepts. It is easy to erase multiple concepts during a generation by either applying steering vectors corresponding to these concepts to the cross-attention output successively or constructing single steering vector corresponding to multiple concepts. In the experiments section, we show results on applying a steering vector constructed for multiple concepts to prevent generation of inappropriate concepts.
Efficiency: Transferring vectors from distilled models. Adversarial Diffusion Distillation (ADD) DBLP:conf/eccv/SauerLBR24 is a fine-tuning approach that allows sampling large-scale foundational image diffusion models in $1$ to $4$ steps, while producing high-quality images, with many methods such as SDXL and Sana having distilled versions (SDXL-Turbo and Sana-Sprint). We observe that steering vectors obtained from the distilled models can successfully be used for steering generations of its corresponding non-distilled variants. More formally, having a pair of prompts, we obtain $ca^{pos\_avg}_{i}$ and $ca^{neg\_avg}_{i}$ from the distilled model using $1$ denoising step. Note that there is no second index $t$ as we use only one denoising iteration, i.e. $T=1$. We then construct steering vectors for the concept $X$ as $ca^{X}_{i}=ca^{pos\_avg}_{i}-ca^{neg\_avg}_{i}$ and then use it to steer non-Turbo variant of the model by using $ca^{X}_{i}$ for each denoising step $1\leqslant j\leqslant T$.
@@ 4 Experiments
Implementation details. For a fair comparison, we report our main quantitative results using StableDiffusion-v1.4 (SD-v1.4) DBLP:conf/cvpr/RombachBLEO22. SD-1.4 model does not have a Turbo version, so for these experiments we use per-step steering vectors computed from the original SD-1.4. We apply steering to all of the CA layers in the model. We set $\beta=2$ for the concept erasure in all experiments. We use 50 prompt pairs for generating steering vectors for concrete concepts (e.g., “Snoopy”), and 196 prompt pairs for generating steering vectors for concrete concepts (e.g., “nudity”). Information about prompts used for generation of the steering vectors is in the supplementary.
We also show effectiveness of CASteer on bigger models, such as SDXL and SANA, and present results on these models in supplementary. For SDXL DBLP:conf/iclr/PodellELBDMPR24 and SANA, we use steering vectors obtained from SDXL-Turbo and SANA-Sprint models, respectively.
Table 1: Quantitative results on nudity removal based on I2P (DBLP:conf/cvpr/SchramowskiBDK23) dataset. Detection of nude body parts is done by Nudenet at a threshold of 0.6. F: Female, M: Male. The best results are highlighted in bold, second-best are underlined.
@@
Ablating (CA) DBLP:conf/iccv/KumariZWS0Z23 298 22 67 7 45 66 180 153 838
FMN zhang2023forgetmenot 155 17 19 2 12 59 117 43 424
ESD-x DBLP:conf/iccv/GandikotaMFB23 101 6 16 10 12 37 77 53 312
SLD-Med DBLP:conf/cvpr/SchramowskiBDK23 39 1 26 3 3 21 72 47 212
UCE DBLP:conf/wacv/GandikotaOBMB24 35 5 11 4 7 29 62 29 182
SA DBLP:conf/nips/HengS23 39 9 4 0 15 32 49 15 163
ESD-u DBLP:conf/iccv/GandikotaMFB23 14 1 8 5 5 24 31 33 121
Receler huang2023receler 13 1 12 9 5 10 26 39 115
MACE DBLP:conf/cvpr/LuWLLK24 16 0 9 7 2 39 19 17 109
RECE DBLP:conf/eccv/GongCWCJ24 8 0 6 4 0 8 23 17 66
CPE (one word) lee2024cpe 11 2 3 2 5 15 13 15 66
CPE (four word) lee2024cpe 6 1 3 2 2 8 8 10 40
AdvUnlearn zhang2024defensive 1 1 0 0 0 13 0 8 23
SAeUron cywinski2025saeuron 4 0 0 1 3 2 1 7 18
@@
Table 2: Quantitative results on inappropriate content removal based on I2P (DBLP:conf/cvpr/SchramowskiBDK23) dataset. Detection of inappropriate content is done by Q16 (schramowski2022can) classifier. The best results are highlighted in bold, second-best are underlined.

### v4

@@ 1 Introduction
Recent advances in diffusion models Ho et al. (2020); Rombach et al. (2022) have revolutionized image Podell et al. (2024) and video generation Girdhar et al. (2024), achieving unprecedented realism. These models operate by gradually adding noise to data during a forward process and then learning to reverse this noise through a series of iterative steps, reconstructing the original data from randomness. By leveraging this denoising process, diffusion models generate high-quality, realistic outputs, making them a powerful tool for creative and generative tasks.
@@
Existing methods for concept erasure in diffusion models remain narrow in scope. LoRA-based fine-tuning Hu et al. (2022) is effective for removing specific objects or styles but struggles with abstract or composite concepts (e.g., nudity, violence, or ideological symbolism), and scales poorly when multiple concepts must be removed, requiring separate adapters or costly retraining. Prompt-based interventions Yoon et al. (2024) offer greater flexibility for abstract harm reduction but lack precision in suppressing concrete attributes, often failing to generalize across concept variations. As a result, existing strategies fall short of delivering reliable, efficient, and broad-spectrum concept erasure.
In this work, we introduce CASteer, a training-free method for controllable concept erasure that leverages the principle of steering to influence hidden representations of diffusion models dynamically. Our method builds on recent findings that deep neural networks encode features into approximately linear subspaces Elhage et al. (2021); Wu et al. (2023). Prior research has shown that intermediate subspaces of diffusion backbones also exhibit this property, with directions that modulate the strength of particular features Kwon et al. (2023); Park et al. (2023); Si et al. (2024); Tumanyan et al. (2023); Li et al. (2024). Yet, these techniques remain limited in scope, often restricted to specific subspaces, requiring training, or offering only coarse control.
Our approach departs from this paradigm. We show that multiple subspaces within diffusion models exhibit linear properties that can be harnessed for precise concept erasure. For each concept of interest, we generate $k$ positive images (where $k\geq 1$) containing the concept and $k$ negative images not containing it, and compute the steering vectors by subtracting the averaged hidden representations of the network across negative images from those of positive ones. During inference, these precomputed vectors are applied directly to the model activations, allowing us to selectively suppress undesirable concepts without retraining or degrading the overall image quality. Experiments demonstrate that CASteer achieves fine-grained erasure of harmful or unwanted concepts (e.g., nudity, violence), while maintaining robustness across a wide range of diffusion models, including SD 1.4, SDXL Podell et al. (2024), Sana Xie et al. (2025), and their distilled variants (e.g., SDXL-Turbo Sauer et al. (2022), Sana-Sprint Chen et al. (2025)).
@@ 2 Related work
Data-driven AI Safety. Ensuring the safety of image and text-to-image generative models hinges on preventing the generation of harmful or unwanted content. Common approaches include curating training data with licensed material Rao (2023); Schuhmann et al. (2022), fine-tuning models to suppress harmful outputs Rombach et al. (2022); Shi et al. (2020), or deploying post-hoc content detectors Bedapudi (2022); Rando et al. (2022). While promising, these strategies face critical limitations: data filtering introduces inherent biases Shi et al. (2020), detectors are computationally efficient but often inaccurate or easily bypassed Gandikota et al. (2023); SmithMano (2022), and model retraining becomes costly when new harmful concepts emerge. Alternative methods leverage text-domain interventions, such as prompt engineering Shi et al. (2020) or negative prompts Miyake et al. (2023); Schramowski et al. (2023). Yet these remain vulnerable to adversarial attacks, lack flexibility and precision as they operate in the discrete space of tokens, and often fail to address the disconnect between text prompts and visual outputs—models can still generate undesired content even when text guidance is “safe”. Our approach instead operates in the joint image-text latent space of diffusion models, enabling more robust and granular control over generated content without relying solely on textual constraints.
@@
Model-driven AI Safety. Current methods Gandikota et al. (2023); Kumari et al. (2023); Heng and Soh (2023); Zhang et al. (2024a); Huang et al. (2024); Lee et al. (2025) erase unwanted concepts by fine-tuning or otherwise optimising models and adapters to shift probability distributions toward null or surrogate tokens, often combined with regularization or generative replay Shin et al. (2017). Other methods, such as Gandikota et al. (2024); Gong et al. (2024), use direct weight editing to remove unwanted concepts. Although effective, these approaches lack precision, inadvertently altering or removing unrelated concepts. Advanced techniques like SPM Lyu et al. (2024) and MACE Lu et al. (2024) improve specificity through LoRA adapters Hu et al. (2022), transport mechanisms, or prompt-guided projections to preserve model integrity. However, while promising for concrete concepts (e.g., Mickey Mouse), they still struggle with abstract concepts (e.g., nudity) and require parameter updates. Another group of methods focuses on interventions into internal mechanisms of generative models. Methods like Prompt-to-Prompt Hertz et al. (2023) enable fine-grained control over text-specified concepts (e.g., amplifying or replacing elements) through interventions to cross-attention maps, yet fail to fully suppress undesired content, particularly when concepts are implicit or absent from prompts. This task-specific specialization limits their utility for safety-critical erasure, where complete removal is required. CASteer bridges this gap, enabling precise, universal concept suppression without relying on textual priors or compromising unrelated model capabilities. Another area of research focuses on removing information about undesired concepts from text embeddings that generative models are conditioned on Yoon et al. (2024); Zhang et al. (2024b); Qiu et al. (2024). However, as these methods operate on a discrete space of token embeddings, their trade-off between the effectiveness of erasure and the preservation of other features is limited. Zhang et al. (2024c) proposes using adversarial training for concept unlearning; however, training this method is computationally intensive. In contrast, CASteer eliminates training entirely, enabling direct, non-invasive concept suppression in the model’s latent space without collateral damage to unrelated features.
@@
Utilizing directions in latent spaces. This area of research focuses on finding interpretable directions in various intermediate spaces of diffusion models Kwon et al. (2023); Park et al. (2023); Si et al. (2024); Tumanyan et al. (2023), which can then be used to control the semantics of generated images. Based on this idea, SDID Li et al. (2024) recently proposed to learn a vector for each given concept, which is then added to the intermediate activation of a bottleneck layer of the diffusion model during inference to provoke the presence of this concept in the generated image. However, this method is highly architecture-specific and fails to deliver precise control over attributes. In our work, we propose a training-free method for constructing interpretable directions in intermediate activation spaces of various diffusion models for more precise control of image generation. SAeUron Cywiński and Deja (2025) utilizes Sparse Autoencoders Olshausen and Field (1997) (SAEs) to find interpretable directions in the activation space of the diffusion model. However, SAEs are unstable, require extensive training, and do not provide initial control over the set of attributes that can be erased. In contrast, CASteer does not require training and provides direct control over the manipulated attributes.
@@ 3.1 Choice of layers to steer
Most modern diffusion models use U-Net or Diffusion Transformers (DiT) Peebles and Xie (2023) as a backbone. They consist of a set of Transformer blocks, each having three main components: cross-attention (CA) layer, self-attention (SA) layer, and MLP layer, all of which contribute to the residual stream of the model. Among those, CA layers are the only place in the model where information from the text prompt goes into the model, guiding text-to-image generation. For every image patch and prompt embedding, each CA layer generates a vector matching the size of the image patch embedding. After summation, these vectors transmit text-prompt information to corresponding image regions Hertz et al. (2023).
As the semantics of the resulting image is mostly determined by the text prompt, we modify the outputs of the CA layers during inference, which results in effective, yet precise, control over the features of the generated image. Thus, CASteer constructs steering vectors for the outputs of every CA layer in the model. In the appendix, we present experiments on applying CASteer to steer outputs of other layers (SA, MLP, and outputs of intermediate layers inside CA blocks).
@@ 3.2 Construction of steering vectors
Figure 1: Main pipeline. (Bottom left, gray background) For computing a steering vector, we prompt diffusion model with two prompts that differ in a desired concept, e.g., “anime style” and save CA outputs at each timestamp $t$ and each CA layer $i$. We average these outputs over image patches and get averaged CA outputs $ca^{pos\_avg}_{it}$ and $ca^{neg\_avg}_{it}$ for each $t$ and $i$. We subtract the latter from the former, getting a steering vector for the layer $i$ and timestamp $t$ $ca^{anime}_{it}$. (Right) For deleting concept $X$ from generation, at each denoising step $t$, we subtract steering vector $ca^{anime}_{it}$ multiplied by intensity $\alpha$ from the CA outputs of the layer $i$.
@@ 3.3 Using steering vectors to control generation
Figure 2: Heatmaps of dot product values between CA outputs of SDXL model (layer 15, denoising step 0) and steering vector for Snoopy concept”. Top: generated images, bottom: heatmaps. Images are generated based on the prompt “A bright photo of a X”, where X $\in$ Snoopy”, Pikachu”, dog”. We see that for the image that contains Snoopy (left), dot product values are high for those image tokens that correspond to the image tokens that actually produce Snoopy. For images that do not contain Snoopy, dot product values are low for all image tokens.
We propose to estimate $\alpha$ for concept deletion by using the dot product between $ca^{X}_{it}$ and corresponding CA output $ca^{out}_{itk}$ ($\langle ca^{X}_{it},ca^{out}_{itk}\rangle$) as an assessment of amount of $X$ that is present in the image part corresponding to $k^{th}$ patch of $ca^{out}_{it}$ (see Fig. 2). As $ca^{X}_{it}$ is normalized, the value of this dot product is the length of the projection of the CA output $ca^{out}_{itk}$ onto the steering vector $ca^{X}_{it}$. As $ca^{X}_{it}$ can be seen as a direction in a linear subspace corresponding to the concept $X$, the length of the projection can be seen as the amount of $X$ that is present in $ca^{out}_{itk}$. That said, for removing information about $X$ from $ca^{out}_{itk}$, we propose to subtract the amount of $ca^{X}_{it}$ proportionate to the dot product between $ca^{X}_{it}$ and $ca^{out}_{itk}$ from $ca^{out\_new}_{itk}$, i.e., define $\alpha=\beta(ca^{X}_{it},ca^{out}_{itk})$. Consequently, Eq. 3 becomes the following:
@@ 3.4 Practical considerations
Steering multiple concepts. It is easy to erase multiple concepts during a generation by either mutually orthogononalizing a set of steeting vectors corresponding to these concepts and applying them to the cross-attention output successively, or constructing single steering vector corresponding to multiple concepts by simply averaging individual steering vectors. In the experiments section, we show results on applying a steering vector constructed for multiple concepts using avergaing approach to prevent generation of inappropriate concepts. Additionally, in the appendix (Sec. M) we show results on erasing multiple concepts by orthogonolizing their corresponding steering vectors.
Efficiency: Transferring vectors from distilled models. Adversarial Diffusion Distillation (ADD) Sauer et al. (2022) is a fine-tuning approach that allows sampling large-scale foundational image diffusion models in $1$ to $4$ steps, while producing high-quality images, with many methods such as SDXL and Sana having distilled versions (SDXL-Turbo and Sana-Sprint). We observe that steering vectors obtained from the distilled models can successfully be used for steering generations of its corresponding non-distilled variants. More formally, having a pair of prompts, we obtain $ca^{pos\_avg}_{i}$ and $ca^{neg\_avg}_{i}$ from the distilled model using $1$ denoising step. Note that there is no second index $t$ as we use only one denoising iteration, i.e. $T=1$. We then construct steering vectors for the concept $X$ as $ca^{X}_{i}=ca^{pos\_avg}_{i}-ca^{neg\_avg}_{i}$ and then use it to steer non-Turbo variant of the model by using $ca^{X}_{i}$ for each denoising step $1\leqslant j\leqslant T$.
@@ 4 Experiments
Implementation details. For a fair comparison, we report our main quantitative results using StableDiffusion-v1.4 (SD-v1.4) Rombach et al. (2022). SD-1.4 model does not have a Turbo version, so for these experiments we use per-step steering vectors computed from the original SD-1.4. We apply steering to all of the CA layers in the model. We set $\beta=2$ for the concept erasure in all experiments. This choice is motivated by the fact that with $\beta=2$, the Eq. 5 becomes a Householder operator (reflection) of the CA activation vector $c$ across the hyperplane orthogonal to the steering vector $s$. This operation preserves $L_{2}$-norm of the vector $c$, thus keeping relative and absolute values of all the information present in $c$ that is orthogonal to $s$ intact after transformation. In appendix (see Sec. H.3) we ablate the choice of $\beta$. We show $\beta<2$ leads to lower level of concept suppression, while still producing high quality images, enabling control on the level of concept erasure. We also show that for SDXL and SANA models, values of $\beta>2$ lead to stronger concept erasure, while also leaving general quality of images high. We use 50 prompt pairs for generating steering vectors for concrete concepts (e.g., “Snoopy”), and 196 prompt pairs for generating steering vectors for concrete concepts (e.g., “nudity”). Information about prompts used for generation of the steering vectors is in the appendix.
We also show effectiveness of CASteer on bigger models, such as SDXL and SANA, and present results on these models in the appendix. For SDXL Podell et al. (2024) and SANA, we use steering vectors obtained from SDXL-Turbo and SANA-Sprint models, respectively.
Table 1: Quantitative results on nudity removal based on I2P Schramowski et al. (2023) dataset. Detection of nude body parts is done by Nudenet at a threshold of 0.6. F: Female, M: Male. The best results are highlighted in bold, second-best are underlined.
@@
DoCo Wu et al. (2025) 162 29 48 63 64 122 168 250 906
Ablating (CA) Kumari et al. (2023) 298 22 67 7 45 66 180 153 838
FMN Zhang et al. (2024a) 155 17 19 2 12 59 117 43 424
ESD-x Gandikota et al. (2023) 101 6 16 10 12 37 77 53 312
SLD-Med Schramowski et al. (2023) 39 1 26 3 3 21 72 47 212
UCE Gandikota et al. (2024) 35 5 11 4 7 29 62 29 182
SA Heng and Soh (2023) 39 9 4 0 15 32 49 15 163
ESD-u Gandikota et al. (2023) 14 1 8 5 5 24 31 33 121
Receler Huang et al. (2024) 13 1 12 9 5 10 26 39 115
MACE Lu et al. (2024) 16 0 9 7 2 39 19 17 109
RECE Gong et al. (2024) 8 0 6 4 0 8 23 17 66
CPE (one word) Lee et al. (2025) 11 2 3 2 5 15 13 15 66
CPE (four word) Lee et al. (2025) 6 1 3 2 2 8 8 10 40
AdvUnlearn Zhang et al. (2024c) 1 1 0 0 0 13 0 8 23
SAeUron Cywiński and Deja (2025) 4 0 0 1 3 2 1 7 18
@@
Table 2: Quantitative results on inappropriate content removal based on I2P Schramowski et al. (2023) dataset. Detection of inappropriate content is done by Q16 Schramowski et al. (2022) classifier. The best results are highlighted in bold, second-best are underlined.